# Tangent spaces of spherical Schubert varieties and a counterexample to the reducedness conjecture

### 18 Mar 2026 v1

@@ Abstract.
Given a simply-connected simple algebraic group $G$, we determine the tangent spaces of the Finkerberg-Mirković Schubert scheme in the affine Grassmannian of $G$ associated to the quasi-miniscule coweight. As a consequence, we exhibit a non-reduced Finkerberg-Mirković Schubert scheme when $G$ is of type $E_{8}$.
@@ 1. Introduction
Let $G$ be a simply-connected simple algebraic group over $\mathbb{C}$ and let $\operatorname{\mathrm{Gr}}_{G}$ denote the affine Grassmannian of $G$. For any dominant coweight $\lambda$, let $\operatorname{\mathrm{Gr}}_{\leq\lambda}$ denote the associated spherical Schubert variety, which is the reduced closure of the $L^{+}G$-orbit associated to $\lambda$. On the other hand, Finkelberg and Mirković proposed in FM a modular description of $\operatorname{\mathrm{Gr}}_{\leq\lambda}$ . It is easy to verify that this moduli space is set-theoretically supported on the Schubert varieties, but it is not at all clear that this moduli space is a reduced scheme. We refer to this moduli space as a Finkelberg-Mirković Schubert scheme (abbreviated as FM Schubert scheme), denoted by $\operatorname{\mathrm{Gr}}^{\rm FM}_{\leq\lambda}$. Since then, it has been a long-standing open question whether FM Schubert schemes are reduced, that is whether they agree with the usual spherical Schubert varieties. This was explicitly formulated as a conjecture in KMW ; see also [KWWY, , Conjecture 2.14, conjecture 2.10], and the discussions in [Zhu, , Remark 2.1.7]. Subsequently, Kamnitzer-Muthiah-Weekes-Yacobi proved that the conjecture holds for type $A$ in KMWY .
@@
Inspired by the work KPZ , we explicitly compute the dimension of the tangent spaces of the spherical Schubert variety $\operatorname{\mathrm{Gr}}_{\leq\theta^{\vee}}$ and the FM Schubert scheme $\operatorname{\mathrm{Gr}}^{\rm FM}_{\leq\theta^{\vee}}$ at the base point $e$, where $\theta^{\vee}$ is the quasi-minisucle coweight, equivalently, the coroot of the highest root $\theta$ of $G$. We show that in Corollary 4.6, except in type $E_{8}$, these two tangent spaces agree and have the same dimension as $G$; in the case of $E_{8}$, however we prove that
@@ 2. Notation and preliminaries
For any $\eta\in X^{*}(T)^{+}$ we write $V_{\eta}$ for the associated irreducible highest-weight representation of $G$ of highest weight $\eta$. We denote the dimension of $\dim V_{\eta}$ by $d_{\eta}$.
@@ 4. Tangent spaces of FM Schubert schemes
over the base point $e\in\operatorname{\mathrm{Gr}}_{G}(\mathbb{C})$. In terms of the Plücker moduli interpretation, $T_{e}\operatorname{\mathrm{Gr}}_{G}$ can be identified with set of those collections $\{L_{\eta,\mathbb{C}[\epsilon]}\}_{\eta\in X^{*}(T)^{+}}$ such that for each $\eta\in X^{*}(T)^{+}$,
@@
We now apply this theorem to the quasi-miniscule coweight $\theta^{\vee}$ of $G$.

### 19 Mar 2026 v2

@@ Abstract.
Given a simply-connected simple algebraic group $G$, we determine the tangent space of the Finkelberg-Mirković Schubert scheme in the affine Grassmannian of $G$ associated to the quasi-minuscule coweight. As a consequence, we exhibit a non-reduced Finkelberg-Mirković Schubert scheme when $G$ is of type $E_{8}$.
2020 Mathematics Subject Classification:
14M15, 22E67
@@ 1. Introduction
Let $G$ be a simply-connected simple algebraic group over $\mathbb{C}$ and let $\operatorname{\mathrm{Gr}}_{G}$ denote the affine Grassmannian of $G$. For any dominant coweight $\lambda$, let $\operatorname{\mathrm{Gr}}_{\leq\lambda}$ denote the associated spherical Schubert variety, which is the reduced closure of the $L^{+}G$-orbit associated to $\lambda$. On the other hand, Finkelberg and Mirković proposed in FM a modular description of $\operatorname{\mathrm{Gr}}_{\leq\lambda}$ . It is easy to verify that this moduli space is set-theoretically supported on the Schubert variety, but it is not at all clear that this moduli space is a reduced scheme. We refer to this moduli space as a Finkelberg-Mirković Schubert scheme (abbreviated as FM Schubert scheme), denoted by $\operatorname{\mathrm{Gr}}^{\rm FM}_{\leq\lambda}$. Since then, it has been a long-standing open question whether FM Schubert schemes are reduced, that is whether they agree with the usual spherical Schubert varieties. This was explicitly formulated as a conjecture in KMW ; see also [KWWY, , Conjecture 2.14, conjecture 2.10], and the discussions in [Zhu, , Remark 2.1.7]. Subsequently, Kamnitzer-Muthiah-Weekes-Yacobi proved that the conjecture holds for type $A$ in KMWY . As explained in KWWY , the validity of this conjecture also has significant applications to the quantization of affine Grassmannian slices and Yangians.
@@
Inspired by the work KPZ , we explicitly compute the dimensions of the tangent spaces of the spherical Schubert variety $\operatorname{\mathrm{Gr}}_{\leq\theta^{\vee}}$ and the FM Schubert scheme $\operatorname{\mathrm{Gr}}^{\rm FM}_{\leq\theta^{\vee}}$ at the base point $e$, where $\theta^{\vee}$ is the quasi-minusucle coweight, equivalently, the coroot of the highest root $\theta$ of $G$. We show that in Corollary 4.6, except in type $E_{8}$, these two tangent spaces agree and have the same dimension as $G$; in the case of $E_{8}$, however we prove that
@@ 2. Notation and preliminaries
For any $\eta\in X^{*}(T)^{+}$ we write $V_{\eta}$ for the associated irreducible highest-weight representation of $G$ of highest weight $\eta$. We denote the dimension of $V_{\eta}$ by $d_{\eta}$.
@@ 4. Tangent spaces of FM Schubert schemes
over the base point $e\in\operatorname{\mathrm{Gr}}_{G}(\mathbb{C})$. In terms of the Plücker moduli interpretation, $T_{e}\operatorname{\mathrm{Gr}}_{G}$ can be identified with the set of those collections $\{L_{\eta,\mathbb{C}[\epsilon]}\}_{\eta\in X^{*}(T)^{+}}$ such that for each $\eta\in X^{*}(T)^{+}$,
@@
We now apply this theorem to the quasi-minuscule coweight $\theta^{\vee}$ of $G$.